# Global regularity for solutions of magnetohydrodynamic equations with large initial data

### v6

@@ 1. Introduction
Then for each $L_{N}\geq 1$ there exist two positive numbers $c=c(N,\nu,\eta,L_{N})$ and $s_{3}=s_{3}(N,L_{N})$ such that
(1.5) $\|w\|_{\infty,Q_{T}}$ $\leq$ $32\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{3}}.$
@@
This combined with Theorem 1.1 implies that a local-in-time strong solution never develops singularity. As a result, it can be extended as a global strong solution.
@@
In the incompressible Navier-Stokes equations, which are (1.1) and (1.3) with $b\equiv 0$, $f_{i}$ is roughly $v_{i}\sum_{j=1}^{N}v_{j}$. As a result, (1.7) becomes
(1.8) $\|v\|_{\infty,Q_{T}}\leq\|v^{(0)}\|_{\infty,\mathbb{R}^{N}}+c\|v\|_{2q,Q_{T}}^{2}.$
Thus, to be able to apply the classical theorem here, one must have $v\in L^{2q}(Q_{T})$ for some $q>N+2$. Of course, this integrability requirement can be weakened due to the work of Serrin, Prodi, and Ladyzenskaja [7]. In spite of that, the gap between what is needed and what is available in terms of integrability is still substantial because we only have
@@
$L_{1}^{j\alpha}\left\|\varphi+\psi\right\|_{\ell,Q_{T}}^{\frac{j\alpha\ell}{\ell-r}}\leq k^{j\alpha}\ \ \mbox{and}\ \ L_{2}^{2j\alpha+1}\|w\|_{r,Q_{T}}^{-(2j\alpha+1)}\|w\|_{q,Q_{T}}^{\frac{(2j\alpha+1)q}{q-L_{N}}}\leq k^{2j\alpha+1}.$
@@
$y_{n+1}$ $\leq$ $\frac{c4^{n}\|w\|_{r,Q_{T}}^{2j\alpha}\|w\|_{q,Q_{T}}^{b}k^{3j\alpha}}{L_{1}^{j\alpha}L_{2}^{2j\alpha+1}\left\|\varphi+\psi\right\|_{\ell,Q_{T}}^{\frac{j\alpha\ell}{\ell-r}}}y_{n}^{1+\alpha}$
@@
(2.21) $b=2-\frac{(2j\alpha+1)q}{q-L_{N}}=\frac{q-2L_{N}}{q-L_{N}}-\frac{2j\alpha q}{q-L_{N}}.$
@@
(2.23) $\frac{2^{r+1}}{k^{r-3j}}\leq\frac{L_{1}^{j}L_{2}^{\frac{2\alpha j+1}{\alpha}}\left\|\varphi+\psi\right\|_{\ell,Q_{T}}^{\frac{j\ell}{\ell-r}}}{c^{\frac{1}{\alpha}}4^{\frac{1}{\alpha^{2}}}\|w\|_{r,Q_{T}}^{2j}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha}}}.$
@@
$+c_{1}4^{-\frac{j}{r-3j}}L_{1}^{-\frac{j}{r-3j}}L_{2}^{-\frac{2\alpha j+1}{\alpha(r-3j)}}\left\|\varphi+\psi\right\|_{\ell,Q_{T}}^{-\beta_{1}}\|w\|_{r,Q_{T}}^{\frac{2j}{r-3j}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha\left(r-3j\right)}},$
@@
(2.25) $c_{1}$ $=$ $2^{\frac{r+1}{r-3j}}c^{\frac{1}{\alpha\left(r-3j\right)}}4^{\frac{1}{\alpha^{2}\left(r-3j\right)}+\frac{j}{r-3j}},$
(2.26) $\beta_{1}$ $=$ $\frac{j\ell}{\left(r-3j\right)(\ell-r)}.$
@@
$+c_{1}4^{-\frac{j}{r-3j}}L_{1}^{-\frac{j}{r-3j}}L_{2}^{-\frac{2\alpha j+1}{\alpha(r-3j)}}\left\|\varphi+\psi\right\|_{\ell,Q_{T}}^{-\beta_{1}}\|w\|_{r,Q_{T}}^{\frac{2j}{r-3j}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha\left(r-3j\right)}}.$
@@
$+2c_{1}L_{2}^{-\frac{2\alpha j+1}{\alpha(r-3j)}}\left\|\varphi+\psi\right\|_{\ell,Q_{T}}^{-\beta_{1}}\|w\|_{r,Q_{T}}^{\frac{2j}{r-3j}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha\left(r-3j\right)}}.$
@@
$+2c_{1}L_{2}^{-\frac{2\alpha j+1}{\alpha(r-3j)}}\left\|w\right\|_{\ell,Q_{T}}^{-\beta_{1}}\|w\|_{r,Q_{T}}^{1+\frac{2j}{r-3j}+\beta_{1}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha\left(r-3j\right)}}.$
@@
$\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+4c_{2}\|w\|_{L_{N},Q_{T}}^{s_{1}}\|w\|_{\ell,Q_{T}}^{-\beta_{1}}\|w\|_{r,Q_{T}}^{1+\frac{2j}{r-3j}+\beta_{1}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha\left(r-3j\right)}},$
@@
(2.30) $c_{2}=c_{1}8^{\frac{2\alpha j+1}{\alpha(r-3j)}},\ \ s_{1}=\frac{L_{N}(2\alpha j+1)}{\alpha(q-L_{N})\left(r-3j\right)}.$
Note from (2.25) that the constant $c_{2}$ here depends on $r,j,q$, and $N$, but not on $\ell$. With this in mind, we will simply denote $4c_{2}$ by $c$ from here on. That is,
(2.31) $\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{1}}\|w\|_{\ell,Q_{T}}^{-\beta_{1}}\|w\|_{r,Q_{T}}^{1+\frac{2j}{r-3j}+\beta_{1}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha\left(r-3j\right)}}.$
@@
Define
(2.32) $M_{q}=\frac{q-2L_{N}}{\alpha q}\ \ \mbox{for $q>\max\{N+2,2L_{N}\}$}.$
Then for each $r>M_{q}$ there exist two positive numbers $c_{3},s_{2}$ such that
(2.33) $\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c_{3}\|w\|_{L_{N},Q_{T}}^{s_{2}}\|w\|_{r,Q_{T}}^{\frac{r}{r-M_{q}}}.$
Moreover, the dependence of $c_{3}$ and $s_{2}$ on $r$ is such that both $\ln c_{3}$ and $s_{2}$ are smooth functions of $r$ over $(M_{q},\infty)$.
@@
(2.34) $\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{1}}\|w\|_{\infty,Q_{T}}^{-\frac{j}{r-3j}}\|w\|_{r,Q_{T}}^{\frac{r}{r-3j}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha\left(r-3j\right)}}.$
Here the constant $c$ remains the same because it is indepenent of $\ell$. Note that (2.29) is still valid. Raise both side of the inequality to the power of $\frac{jq}{(r-3j)(q-L_{N})}$ to deduce
$\|w\|_{q,Q_{T}}^{\frac{jq}{(r-3j)(q-L_{N})}}\leq\|w\|_{\infty,Q_{T}}^{\frac{j}{r-3j}}\|w\|_{L_{N},Q_{T}}^{\frac{jL_{N}}{(r-3j)(q-L_{N})}}.$
@@
(2.35) $\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{2}}\|w\|_{r,Q_{T}}^{\frac{r}{r-3j}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha\left(r-3j\right)}-\frac{jq}{(r-3j)(q-L_{N})}}.$
@@
(2.36) $s_{2}=s_{1}+\frac{jL_{N}}{(r-3j)(q-L_{N})}.$
We pick our parameters in such a way that
(2.37) $\frac{b}{\alpha\left(r-3j\right)}-\frac{jq}{(r-3j)(q-L_{N})}=0.$
@@
$\frac{q-2L_{N}-3\alpha jq}{\alpha(r-3j)(q-L_{N})}=0,$
@@
$j=\frac{q-2L_{N}}{3\alpha q}=\frac{M_{q}}{3}.$
We arrive at (2.33) by inserting the preceding choice of $j$ into (2.35). We easily check from (2.25), (2.30), and (2.36) that
(2.38) $s_{2}$ $=$ $\frac{L_{N}(2\alpha M_{q}+3)}{3\alpha(q-L_{N})\left(r-M_{q}\right)}+\frac{M_{q}L_{N}}{3(r-M_{q})(q-L_{N})},$
(2.39) $c_{3}$ $=$ $2^{\frac{r+1}{r-M_{q}}+2}c^{\frac{1}{\alpha\left(r-M_{q}\right)}}4^{\frac{1}{\alpha^{2}\left(r-M_{q}\right)}+\frac{M_{q}}{3(r-M_{q})}}8^{\frac{2\alpha M_{q}+3}{3\alpha(r-M_{q})}}.$
Note that the constant $c$ on the right-hand side of (2.39) is the same as the one in (2.25), and hence it depends only on $q$. The proof is complete. ∎
@@
$\lim_{q\rightarrow\infty}M_{q}=\frac{N+2}{2}.$
@@
This implies that we can take $r=L_{N}$ in (2.33). That is to say, $w\in L^{\infty}(Q_{T})$ whenever (2.40) holds.
@@
$L_{N}\leq\frac{N+2}{2}.$
With this in mind, we can check from (2.18) that
$M_{q}=\frac{(N+2)(q-2L_{N})}{2(q-N-2)}\geq\frac{N+2}{2}.$
It turns out that we can improve (2.33) by making better choice of the parameters in (2.31). To this end, we consider the function
$f(r)=\ln\left(\int_{Q_{T}}w^{r}dxdt\right)\ \ \mbox{for $r>M_{q}$}.$
@@
$=$ $e^{\left(\frac{f(r)}{r}\right)^{\prime}}$
$=$ $\|w\|_{r,Q_{T}}^{-\frac{1}{r}}e^{\frac{\int_{Q_{T}}w^{r}\ln wdxdt}{r\int_{Q_{T}}w^{r}dxdt}}.$
Remember that the constant $c$ in (2.31) does not depend $\ell$. With this in mind, we substitute (2.26) into (2.31) and take $\ell\rightarrow r$ in the resulting inequality, thereby obtaining
@@
$+c\|w\|_{L_{N},Q_{T}}^{s_{1}}\left[\lim_{\ell\rightarrow r}\left(\frac{\|w\|_{\ell,Q_{T}}}{\|w\|_{r,Q_{T}}}\right)^{\frac{1}{\ell-r}}\right]^{-\frac{j\ell}{r-3j}}\|w\|_{r,Q_{T}}^{\frac{r-j}{r-3j}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha(r-3j)}}$
@@
$+c\|w\|_{L_{N},Q_{T}}^{s_{1}}\left[\|w\|_{r,Q_{T}}^{-1}e^{\frac{\int_{Q_{T}}w^{r}\ln wdxdt}{\int_{Q_{T}}w^{r}dxdt}}\right]^{-\frac{j}{r-3j}}\|w\|_{r,Q_{T}}^{\frac{r-j}{r-3j}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha(r-3j)}}$
$=$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{1}}e^{\frac{-jf^{\prime}(r)}{\left(r-3j\right)}}\|w\|_{r,Q_{T}}^{\frac{r}{r-3j}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha(r-3j)}}.$

### v7

@@ 1. Introduction
Then for each $L_{N}\geq 1$ there exist two positive numbers $c=c(N,\nu,\eta,L_{N})$ and $s_{5}=s_{5}(N,L_{N})$ such that
(1.5) $\|w\|_{\infty,Q_{T}}$ $\leq$ $32\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{5}}.$
@@
This combined with Theorem 1.1 implies that a local-in-time strong solution never develops singularity. As a result, it can be extended as a global strong solution. If $b\equiv 0$, Theorem 1.1 implies a positive answer to the Navier-Stokes millennium problem [5, 11].
@@
In the incompressible Navier-Stokes equations, which are (1.1) and (1.3) with $b\equiv 0$, $f_{i}$ is roughly $v_{i}\sum_{j=1}^{N}v_{j}+p$. As a result, (1.7) becomes
(1.8) $\|v\|_{\infty,Q_{T}}\leq\|v^{(0)}\|_{\infty,\mathbb{R}^{N}}+c\left(\|v\|_{2q,Q_{T}}^{2}+\|p\|_{q,Q_{T}}\right)\leq\|v^{(0)}\|_{\infty,\mathbb{R}^{N}}+c\|v\|_{2q,Q_{T}}^{2}.$
The last step is due to (2.15) below. Thus, to be able to apply the classical theorem here, one must have $v\in L^{2q}(Q_{T})$ for some $q>N+2$. Of course, this integrability requirement can be weakened due to the work of Serrin, Prodi, and Ladyzenskaja [8]. In spite of that, the gap between what is needed and what is available in terms of integrability is still substantial because we only have
@@
$L_{1}^{j\alpha}\left\|\varphi+\psi\right\|_{\ell,Q_{T}}^{\frac{j\alpha\ell}{\ell-r}}\leq k^{j\alpha}\ \ \mbox{and}\ \ L_{2}^{j\alpha+1}\|w\|_{r,Q_{T}}^{-(j\alpha+1)}\|w\|_{q,Q_{T}}^{\frac{(j\alpha+1)q}{q-L_{N}}}\leq k^{j\alpha+1}.$
@@
$y_{n+1}$ $\leq$ $\frac{c4^{n}\|w\|_{r,Q_{T}}^{j\alpha}\|w\|_{q,Q_{T}}^{b}k^{2j\alpha}}{L_{1}^{j\alpha}L_{2}^{j\alpha+1}\left\|\varphi+\psi\right\|_{\ell,Q_{T}}^{\frac{j\alpha\ell}{\ell-r}}}y_{n}^{1+\alpha}$
@@
(2.21) $b=2-\frac{(j\alpha+1)q}{q-L_{N}}=\frac{q-2L_{N}}{q-L_{N}}-\frac{j\alpha q}{q-L_{N}}.$
@@
(2.23) $\frac{2^{r+1}}{k^{r-2j}}\leq\frac{L_{1}^{j}L_{2}^{\frac{\alpha j+1}{\alpha}}\left\|\varphi+\psi\right\|_{\ell,Q_{T}}^{\frac{j\ell}{\ell-r}}}{c^{\frac{1}{\alpha}}4^{\frac{1}{\alpha^{2}}}\|w\|_{r,Q_{T}}^{j}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha}}}.$
@@
$+c_{1}4^{-\frac{j}{r-2j}}L_{1}^{-\frac{j}{r-2j}}L_{2}^{-\frac{\alpha j+1}{\alpha(r-2j)}}\left\|\varphi+\psi\right\|_{\ell,Q_{T}}^{-\beta_{1}}\|w\|_{r,Q_{T}}^{\frac{j}{r-2j}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha\left(r-2j\right)}},$
@@
(2.25) $c_{1}$ $=$ $2^{\frac{r+1}{r-2j}}c^{\frac{1}{\alpha\left(r-2j\right)}}4^{\frac{1}{\alpha^{2}\left(r-2j\right)}+\frac{j}{r-2j}},$
(2.26) $\beta_{1}$ $=$ $\frac{j\ell}{\left(r-2j\right)(\ell-r)}.$
@@
$+c_{1}4^{-\frac{j}{r-2j}}L_{1}^{-\frac{j}{r-2j}}L_{2}^{-\frac{\alpha j+1}{\alpha(r-2j)}}\left\|\varphi+\psi\right\|_{\ell,Q_{T}}^{-\beta_{1}}\|w\|_{r,Q_{T}}^{\frac{j}{r-2j}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha\left(r-2j\right)}}.$
@@
$+2c_{1}L_{2}^{-\frac{\alpha j+1}{\alpha(r-2j)}}\left\|\varphi+\psi\right\|_{\ell,Q_{T}}^{-\beta_{1}}\|w\|_{r,Q_{T}}^{\frac{j}{r-2j}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha\left(r-2j\right)}}.$
@@
$+2c_{1}L_{2}^{-\frac{\alpha j+1}{\alpha(r-2j)}}\left\|w\right\|_{\ell,Q_{T}}^{-\beta_{1}}\|w\|_{r,Q_{T}}^{1+\frac{j}{r-2j}+\beta_{1}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha\left(r-2j\right)}}.$
@@
(2.30) $\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c_{2}\|w\|_{L_{N},Q_{T}}^{s_{1}}\|w\|_{\ell,Q_{T}}^{-\beta_{1}}\|w\|_{r,Q_{T}}^{1+\frac{j}{r-2j}+\beta_{1}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha\left(r-2j\right)}},$
@@
(2.31) $c_{2}=4c_{1}8^{\frac{\alpha j+1}{\alpha(r-2j)}}=c^{\frac{1}{\alpha\left(r-2j\right)}}4^{\frac{1}{\alpha^{2}\left(r-2j\right)}+\frac{r-j}{r-2j}+\frac{2(\alpha j+1)}{\alpha(r-2j)}+\frac{r+1}{2(r-2j)}},\ \ s_{1}=\frac{L_{N}(\alpha j+1)}{\alpha(q-L_{N})\left(r-2j\right)}.$
In summary, we have shown:
@@
For each $q$ satisfying (2.16) there is a constant $c=c(N,q,\nu,\eta)$ such that (2.30) holds for each
$j>0,\ \ r>2j,\ \ \mbox{and}\ \ \ell>r.$
We wish to transform the last three different norms in (2.30) into a single one. Our first attempt yields the following result.
Claim 2.3.
Let $q$ be given as in (2.16). We further assume
(2.32) $q>2L_{N}.$
Then for each
(2.33) $r>M_{q}\equiv\frac{q-2L_{N}}{\alpha q},$
there exist two positive numbers $c_{3},s_{0}$ such that
(2.34) $\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c_{3}\|w\|_{L_{N},Q_{T}}^{s_{0}}\|w\|_{r,Q_{T}}^{\frac{r}{r-M_{q}}}.$
Moreover, the dependence of $c_{3}$ and $s_{0}$ on $r$ is such that both $\ln c_{3}$ and $s_{2}$ are smooth functions of $r$ over $(M_{q},\infty)$.
@@
(2.35) $\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c_{2}\|w\|_{L_{N},Q_{T}}^{s_{1}}\|w\|_{\infty,Q_{T}}^{-\frac{j}{r-2j}}\|w\|_{r,Q_{T}}^{\frac{r}{r-2j}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha\left(r-2j\right)}}.$
Here the constants $c_{2}$ and $s_{1}$ remain the same as those in (2.30) because they are independent of $\ell$. Under (2.32), inequality (2.29) remains valid. Raise both side of the inequality to the power of $\frac{jq}{(r-2j)(q-L_{N})}$ to deduce
$\|w\|_{q,Q_{T}}^{\frac{jq}{(r-2j)(q-L_{N})}}\leq\|w\|_{\infty,Q_{T}}^{\frac{j}{r-2j}}\|w\|_{L_{N},Q_{T}}^{\frac{jL_{N}}{(r-2j)(q-L_{N})}}.$
@@
(2.36) $\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c_{2}\|w\|_{L_{N},Q_{T}}^{s_{2}}\|w\|_{r,Q_{T}}^{\frac{r}{r-2j}}\|w\|_{q,Q_{T}}^{\frac{b}{\alpha\left(r-2j\right)}-\frac{jq}{(r-2j)(q-L_{N})}}.$
@@
(2.37) $s_{2}=s_{1}+\frac{jL_{N}}{(r-2j)(q-L_{N})}.$
We pick $j$ so that
(2.38) $\frac{b}{\alpha\left(r-2j\right)}-\frac{jq}{(r-2j)(q-L_{N})}=0.$
@@
$\frac{q-2L_{N}-2\alpha jq}{\alpha(r-2j)(q-L_{N})}=0,$
@@
$j=\frac{q-2L_{N}}{2\alpha q}=\frac{M_{q}}{2}.$
We arrive at (2.34) by inserting the preceding choice of $j$ into (2.36). In view of (2.31) and (2.37), we have
$s_{0}$ $=$ $s_{2}|_{j=\frac{M_{q}}{2}}=\frac{L_{N}(\alpha M_{q}+2)}{2\alpha(q-L_{N})\left(r-M_{q}\right)}+\frac{M_{q}L_{N}}{2(r-M_{q})(q-L_{N})},$
(2.39) $c_{3}$ $=$ $c_{2}|_{j=\frac{M_{q}}{2}}=c^{\frac{1}{\alpha\left(r-M_{q}\right)}}4^{\frac{1}{\alpha^{2}\left(r-M_{q}\right)}+\frac{2r-M_{q}}{2(r-M_{q})}+\frac{\alpha M_{q}+2}{\alpha(r-M_{q})}+\frac{r+1}{2(r-M_{q})}}.$
Note that the constant $c$ on the right-hand side of (2.39) is the same as the one in (2.25), and hence it depends on $N,q,\nu,\eta$ only. The proof is complete. ∎
@@
(2.40) $\lim_{q\rightarrow\infty}M_{q}=\frac{N+2}{2}.$
@@
This implies that we can take $r=L_{N}$ in (2.34). That is to say, $w\in L^{\infty}(Q_{T})$ whenever $w\in L^{L_{N}}(Q_{T})$ with $L_{N}$ satisfying (2.41).
@@
(2.42) $L_{N}<\frac{N+2}{2}.$
As a result, we have (2.32). That is to say, (2.34) is true under the assumptions (2.16), (2.42), and (2.33). Moreover, plug (2.18) into the definition of $M_{q}$ in (2.33) to derive
(2.43) $M_{q}=\frac{(N+2)(q-2L_{N})}{2(q-N-2)}>\frac{N+2}{2}>L_{N}.$
We can also represent $b$ in (2.21) as
$b=\frac{\alpha q(M_{q}-j)}{q-L_{N}}.$
Plugging this and (2.26) into (2.30), we arrive at
(2.44) $\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}$
$+c_{2}\|w\|_{L_{N},Q_{T}}^{s_{1}}\|w\|_{\ell,Q_{T}}^{-\frac{j\ell}{(r-2j)(\ell-r)}}\|w\|_{r,Q_{T}}^{1+\frac{j}{r-2j}+\frac{j\ell}{(r-2j)(\ell-r)}}\|w\|_{q,Q_{T}}^{\frac{q(M_{q}-j)}{(q-L_{N})\left(r-2j\right)}}.$
We wish to improve (2.34) by making better choice of the parameters in (2.44). What motivates us to do so is the observation that the sum of the last three exponents in (2.44) is given by
$1+\frac{j}{r-2j}+\frac{q(M_{q}-j)}{(q-L_{N})\left(r-2j\right)}=1+\frac{-L_{N}j+M_{q}q}{(r-2j)(q-L_{N})}.$
Obviously, the total sum can be made smaller than 1 if we choose $j$ suitably large. This suggests that one should try to transform the last three norms in (2.44) into a single one in such a way that the total sum is largely preserved. To this end, we consider the function
$f(r)=\ln\left(\int_{Q_{T}}w^{r}dxdt\right)\ \ \mbox{for $r\geq L_{N}$}.$
@@
That is, $f(r)$ is convex over $(L_{N},\infty)$. We will exploit this property extensively. We may also assume that
$f^{\prime}(r)>0\ \ \mbox{for each $r>M_{q}$.}$
Indeed, suppose that the above is not true, i.e., there is a $r>M_{q}$ such that
(2.47) $f^{\prime}(r)\leq 0.$
By (2.45), we have
$\int_{\{w\geq 1\}}w^{r}\ln wdxdt$ $\leq$ $-\int_{\{w<1\}}w^{r}\ln wdxdt$
$\leq$ $\max_{s\in[0,1]}|s^{r-\frac{N+2}{N}}\ln s|\int_{\{w<1\}}w^{\frac{N+2}{2}}dxdt\leq c.$
The last step here is due to (1.6). For each $L>1$ we have
$\int_{Q_{T}}w^{r}dxdt$ $=$ $\int_{\{w\geq L\}}w^{r}dxdt+\int_{\{w<L\}}w^{r}dxdt$
$\leq$ $\frac{1}{\ln L}\int_{\{w\geq L\}}w^{r}\ln wdxdt+L^{r-\frac{N+2}{N}}\int_{\{w<L\}}w^{\frac{N+2}{N}}dxdt\leq c.$
This combined with (2.34) implies our theorem. That is to say, were (2.47) true for some $r>M_{q}$, we would have nothing more to prove.
@@
$=$ $e^{\left(\frac{f(r)}{r}\right)^{\prime}}=\|w\|_{r,Q_{T}}^{-\frac{1}{r}}e^{\frac{\int_{Q_{T}}w^{r}\ln wdxdt}{r\int_{Q_{T}}w^{r}dxdt}}.$
Remember that the constant $c_{2}$ in (2.44) does not depend $\ell$. For simplicity, from here on we will just denote it by $c$. With this in mind, we take $\ell\rightarrow r$ in (2.44), thereby obtaining
@@
$+c\|w\|_{L_{N},Q_{T}}^{s_{1}}\left[\lim_{\ell\rightarrow r}\left(\frac{\|w\|_{\ell,Q_{T}}}{\|w\|_{r,Q_{T}}}\right)^{\frac{1}{\ell-r}}\right]^{-\frac{j\ell}{r-2j}}\|w\|_{r,Q_{T}}^{\frac{r-j}{r-2j}}\|w\|_{q,Q_{T}}^{\frac{q(M_{q}-j)}{(q-L_{N})(r-2j)}}$
@@
$+c\|w\|_{L_{N},Q_{T}}^{s_{1}}\left[\|w\|_{r,Q_{T}}^{-1}e^{\frac{\int_{Q_{T}}w^{r}\ln wdxdt}{\int_{Q_{T}}w^{r}dxdt}}\right]^{-\frac{j}{r-2j}}\|w\|_{r,Q_{T}}^{\frac{r-j}{r-2j}}\|w\|_{q,Q_{T}}^{\frac{q(M_{q}-j)}{(q-L_{N})(r-2j)}}$
$=$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{1}}e^{\frac{-jf^{\prime}(r)}{\left(r-2j\right)}}\|w\|_{r,Q_{T}}^{\frac{r}{r-2j}}\|w\|_{q,Q_{T}}^{\frac{q(M_{q}-j)}{(q-L_{N})(r-2j)}}.$
$=$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{1}}e^{\frac{-jf^{\prime}(r)}{\left(r-2j\right)}+\frac{f(r)}{r-2j}}\|w\|_{q,Q_{T}}^{\frac{q(M_{q}-j)}{(q-L_{N})(r-2j)}}.$
To continue, we refine our selection of parameters further. In view of (2.40), we may pick
(2.49) $q>M_{q}.$
Subsequently, take
(2.50) $r>q.$
Introduce two new parameters
$A\in(0,\infty)\ \ \mbox{and}\ \ \xi\in(q,r).$
Consider the initial value problem
(2.53) $h^{\prime}(s)+\frac{h(s)}{s-M_{q}}$ $=$ $\left\{\begin{array}[]{ll}1&\mbox{if $s\in(q,\xi]$},\\ -A&\mbox{if $s\in(\xi,\infty)$,}\end{array}\right$.
(2.54) $h(q)$ $=$ $0.$
Solve this to obtain
$h(s)=\left\{\begin{array}[]{ll}\frac{s-M_{q}}{2}-\frac{(q-M_{q})^{2}}{2(s-M_{q})}&\mbox{if $s\in[q,\xi]$},\\ -\frac{A(s-M_{q})}{2}+\frac{(A+1)(\xi-M_{q})^{2}-(q-M_{q})^{2}}{2(s-M_{q})}&\mbox{if $s\in(\xi,\infty]$.}\end{array}\right$.
We can easily verify that $h(s)$ is Lipschitz on $[q,r]$. Furthermore, it is increasing on $(q,\xi)$ and decreasing on $(\xi,r)$. Hence,
(2.55) $h(s)>0\ \ \mbox{on $(q,r]$ whenever $h(r)>0$.}$
To verify the latter, we have
(2.56) $h(r)$ $=$ $-\frac{A(r-M_{q})}{2}+\frac{(A+1)(\xi-M_{q})^{2}-(q-M_{q})^{2}}{2(r-M_{q})}$
$=$ $\frac{-A\left[(r-M_{q})^{2}-(\xi-M_{q})^{2}\right]+(\xi-M_{q})^{2}-(q-M_{q})^{2}}{2(r-M_{q})}$
$=$ $\frac{-A(r-\xi)(r+\xi-2M_{q})+(\xi-q)(\xi+q-2M_{q})}{2(r-M_{q})}.$
Hence, to achieve the first inequality in (2.55), it is enough for us to take
(2.57) $A<\frac{(\xi-q)(\xi+q-2M_{q})}{(r-\xi)(r+\xi-2M_{q})}.$
Note from (2.54) that
(2.58) $\frac{f(r)}{r-2j}$ $=$ $\frac{h(r)f(r)}{h(r)(r-2j)}$
$=$ $\frac{1}{h(r)(r-2j)}\left[h(q)f(q)+\int_{q}^{r}\left(h(s)f(s)\right)^{\prime}ds\right]$
$=$ $\frac{1}{h(r)(r-2j)}\left[\int_{q}^{r}h(s)f^{\prime}(s)ds+\int_{q}^{r}h^{\prime}(s)f(s)ds\right].$
To estimate $f^{\prime}(s)$, we appeal to (2.34). But we will use a slightly different version of it. Without any loss of generality, we may assume that the second term in (2.34) is less that or equal to the third term there, i.e.,
$16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}\leq c_{3}\|w\|_{L_{N},Q_{T}}^{s_{0}}\|w\|_{s,Q_{T}}^{\frac{s}{s-M_{q}}}\ \ \mbox{for each $s>M_{q}$}.$
(We have changed $r$ in (2.34) to $s$.) Were this not true, we would have nothing more to prove. Consequently, (2.34) is reduced to
$\|w\|_{\infty,Q_{T}}$ $\leq$ $2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{0}}\|w\|_{s,Q_{T}}^{\frac{s}{s-M_{q}}}\ \ \mbox{for $s\in(M_{q},\infty)$}.$
This together with (2.45) implies that
$f^{\prime}(s)\leq\ln\left(\|w\|_{\infty,Q_{T}}\right)\leq\ln\left(2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{0}}\right)+\frac{f(s)}{s-M_{q}},\ \ s>M_{q}.$
In view of (2.49), (2.55), and (2.53), we have
(2.59) $\int_{q}^{r}h(s)f^{\prime}(s)ds+\int_{q}^{r}h^{\prime}(s)f(s)ds$
$\leq$ $\ln\left(2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{0}}\right)\int_{q}^{r}h(s)ds+\int_{q}^{r}\left(\frac{h(s)}{s-M_{q}}+h^{\prime}(s)\right)f(s)ds$
$=$ $\ln\left(2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{0}}\right)\int_{q}^{r}h(s)ds+\int_{q}^{\xi}f(s)ds-A\int_{\xi}^{r}f(s)ds.$